OpenStudy (anonymous):

Write out the steps to solve the equation 3n = 64 – 5n please do step by step plz and thx medal goes to the 1st right person

OpenStudy (anonymous):

u 5 senetences if u can

OpenStudy (anonymous):

idk if this may be right but this is how i did it. 3n=64-5n 3-5n=-2 -2n there 64/=-2 is N=-32 hope it's what you looking for.